PLEASE ANSWER ALL PARTS :)

17 (3.38) NOTE: The numerical values in this problem have been modified for testing purposes. In its Fuel Economy Guide for 2014 model vehicles, the Environmental Protection Agency gives data on 1160 vehicles. There are a number of high outliers, mainly hybrid gas-electric vehicles. If we ignore the vehicles identified as outliers, however, the combined city and highway gas mileage of the other 1134 vehicles is approximately Normal with mean 22.9 miles per gallon (mpg) and standard deviation 5.2 mpg The quintiles of any distribution are the values with cumulative proportions O 20 0 40 Ο.60 and (Use software) The first quintile is 0.01 mpg) mpg The second quintile is (0.01 mpg) mpg The third quintile is (0.01 mpg) mpg The fourth quintile is 0.01 mpg) mpg 80 what are the quintiles of the distribution of gas mileage
Answer #1

The quintiles of the distribution of gas mileage (First, second, third and fourth quintiles) can be obtained by using the ‘standard normal procedure’ and the ‘z-score’

a)

first quintile

P(Z<=z) = 0.20

Z0.20 = -0.84 [excel formula =NORMSINV(0.20)]

Z=(X-µ)/σ

µ=22.9

σ=5.2

x=Zσ+µ = -0.84*5.2+22.9=18.532

b)

second quintile

Z0.40=-0.26

x=Zσ+µ = -0.26*5.2+22.9=21.548

c)

third quintile

Z0.60 = 0.25

x=Zσ+µ = 0.25*5.2+22.9=24.2

d)

fourth quintile

Z.80 = 0.84

x=Zσ+µ = 0.84*5.2+22.9=27.268